# Toward Vu’s conjecture

### v1

@@ 5. An exceptional variant of a concentration inequality
Thus, applying Lemma 2.3 with $\beta=1$, $D=\ell$, $\Omega^{*}=\emptyset$, and $\tau=\ell/\log^{5}\ell$, we have
$\Pr\left(\left|\ell^{\prime}(v)-\mathbb{E}[\ell^{\prime}(v)]\right|>\frac{\ell}{\log^{5}\ell}\right)$ $=\Pr\left(\left|R_{v}-\mathbb{E}[R_{v}]\right|>\frac{\ell}{\log^{5}\ell}\right)$
$<6\exp\left(-\frac{\ell^{2}\log^{-10}\ell}{24\ell}\right)=O(\exp(-\log^{2}d)),$
@@
Let $H$ denote the set of vertices $w\in V(G)$ for which $c\in L(w)$ and $w$ has at least $|S|/\log^{13}d$ neighbors in $S$. We first estimate $|H|$. Consider the graph $H^{\prime}$ on $H\cup S$ consisting of all edges with an endpoint in $H$ and an endpoint in $S$. As $(G,L)$ is a $(d,\ell,s,\eta)$-graph list pair, $H^{\prime}$ is $K_{s,t}$-free, where $t:=d/\log^{16s}d+1$. In particular, there is no $Y\subseteq S$ of size $s$ and $X\subseteq H$ of size $t$ such that $X\cap Y=\emptyset$ and $[X,Y]$ forms a $K_{s,t}$. Applying Corollary 2.2 with $n=|S|$ and $m=|H|$, we have
@@
where this follows as $d_{L}(v,c)\leq d$ and $\eta d<\ell$. By a Chernoff bound (Lemma 2.4), we have that
@@
$\frac{|S|}{8\log^{5}|S|}>\frac{d}{12\log^{16s+5}d}>12\sqrt{6\pi|S|\log^{16}d}+2|S|\Pr(\Omega^{*}).$
Thus, we apply Lemma 2.3 with $\beta=3\log^{16}d$, $D=2|S|$ and $\tau=\frac{|S|}{8\log^{5}|S|}$ to get
$\Pr\left(|S_{\mu}|-\mathbb{E}|S_{\mu}|\geq\frac{|S|}{8\log^{5}|S|}\right)\leq 6\exp\left(-\frac{|S|^{2}8^{-2}\log^{-10}|S|}{24(2|S|)\left(3\log^{16}d\right)}\right)+\Pr(\Omega^{*})$
$\leq$ $6\exp\left(\frac{-|S|}{(10^{5}\log^{10}|S|)\,(\log^{16}d)}\right)+O(\exp(-\log^{2}d))=O(\exp(-\log^{2}d)).$
@@
$\frac{|S|}{8\log^{5}|S|}>120\frac{|S|}{\log^{13/2}d}+2\Pr(\Omega^{*})|S|\ >12\sqrt{\pi\frac{2|S|}{\log^{13}d}3|S|}+2\Pr(\Omega^{*})|S|.$
Hence, applying Lemma 2.3 with $\beta=2|S|/\log^{13}d$, $D=3|S|$ and $\tau=\frac{|S|}{8\log^{5}|S|}$, we obtain
@@
Now, we estimate the probability of each bad event. For each $v\in V(G)$, by Lemmas 6.1 and 6.2,
@@
Before we proceed with the proofs, let us record a few inequalities. First, as $d$ is sufficiently large, by Lemmas 2.6 and 7.2, we have
@@
The proof is by induction on $i$. Let $i\geq 1$, and let $r_{i}\coloneqq d_{i}/\ell_{i}$ and $\hat{r}_{i}\coloneqq\hat{d}_{i}/\hat{\ell}_{i}$, where $\hat{d}_{i}$ and $\hat{\ell}_{i}$ are defined as in Lemma 7.3. Now we assume that the desired bound holds for $\ell_{1},\ldots,\ell_{i}$ and consider $\ell_{i+1}$. By the induction hypothesis, we have $\ell_{j}\geq d^{\varepsilon/15}$ for $j\leq i$. Furthermore, as we assume $d_{j}\geq\ell_{j}/8$ and $d$ is sufficiently large, we have $d_{j}\geq d^{\varepsilon/20}$. In particular, we may apply Lemmas 7.2 and 7.3. The following claim will assist with our proof:
@@
Note that as we assume $i<i^{\star}$, we have $\log\ell_{i}\geq\frac{\varepsilon}{15}\log d$ and $\log d_{i}\geq\frac{\varepsilon}{20}\log d$. Therefore, $\eta\gg\log^{-3}\ell_{i}$ and $\eta\gg\log^{-3}d_{i}$, and we can apply Lemmas 7.2 and 7.3.

### v2

@@ 4. The Wasteful Coloring Procedure
Thus, applying Lemma 2.3 with $\beta=r=1$, $\Omega^{*}=\emptyset$, $n=\ell$, and $\tau=\ell/\log^{5}\ell$, we have
$\Pr\left(\left|\ell^{\prime}(v)-\mathbb{E}[\ell^{\prime}(v)]\right|>\frac{\ell}{\log^{5}\ell}\right)$ $=\Pr\left(\left|R-\mathbb{E}[R]\right|>\frac{\ell}{\log^{5}\ell}\right)$
$<4\exp\left(-\frac{\ell^{2}\log^{-10}\ell}{40\ell}\right)=O(\exp(-\log^{2}d)),$
@@
Let $H$ denote the set of vertices $w\in V(G)$ for which $c\in L(w)$ and $w$ has at least $|S|/\log^{13}d$ neighbors in $S$. We first estimate $|H|$. Consider the graph $H^{\prime}$ on $H\cup S$ consisting of all edges with an endpoint in $H$ and an endpoint in $S$. As $(G,L)$ is a $(d,\ell,s,\eta)$-graph list pair, $H^{\prime}$ is $K_{s,t}$-free, where $t:=(d/\log^{16s}d)+1$. In particular, there is no $K_{s,t}$ with parts $X,Y$ where $Y\subseteq S$ is of size $s$ and $X\subseteq H$ is of size $t$. Applying Corollary 2.2 with $n=|S|$ and $m=|H|$, we have
@@
where this follows as $d_{L}(v,c)\leq d$ and $\eta d<\ell$. Using the Chernoff bound (Lemma 2.4) with $\mu=1$ (since $\mathbb{E}[X_{c^{\prime}}]<1$) and $\delta=2\log^{16}d-1$, we have that
@@
$\frac{|S|}{8\log^{5}|S|}>\frac{d}{12\log^{16s+5}d}>96\sqrt{(2)(3\log^{16}d)|S|}+128(2)(3\log^{16}d)+8|S|\Pr(\Omega^{*}).$
Thus, we apply Lemma 2.3 with $\beta=3\log^{16}d$, $r=2$, $n=|S|$, and $\tau=\frac{|S|}{8\log^{5}|S|}$ to get
$\Pr\left(|S_{\mu}|-\mathbb{E}|S_{\mu}|\geq\frac{|S|}{8\log^{5}|S|}\right)\leq 4\exp\left(-\frac{|S|^{2}8^{-2}\log^{-10}|S|}{40\left(3\log^{16}d\right)(2)|S|}\right)+4\Pr(\Omega^{*})$
$\leq$ $4\exp\left(\frac{-|S|}{(10^{6}\log^{10}|S|)\,(\log^{16}d)}\right)+O(\exp(-\log^{2}d))=O(\exp(-\log^{2}d)).$
@@
$\frac{|S|}{8\log^{5}|S|}>300\frac{|S|}{\log^{13/2}d}+8\Pr(\Omega^{*})|S|\ >96\sqrt{\frac{3\cdot 2|S|^{2}}{\log^{13}d}}+128(3)\frac{2|S|}{\log^{13}d}+8\Pr(\Omega^{*})|S|.$
Hence, applying Lemma 2.3 with $\beta=2|S|/\log^{13}d$, $r=3$, $n=|S|$, and $\tau=\frac{|S|}{8\log^{5}|S|}$, we obtain
@@
Now, we estimate the probability of each bad event. For each $v\in V(G)$, by Lemmas 5.1 and 5.2,
@@
Before we proceed with the proofs, let us record a few inequalities. First, as $d$ is sufficiently large, by Lemmas 2.6 and 6.2, we have
@@
The proof is by induction on $i$. Let $i\geq 1$, and let $r_{i}\coloneqq d_{i}/\ell_{i}$ and $\hat{r}_{i}\coloneqq\hat{d}_{i}/\hat{\ell}_{i}$, where $\hat{d}_{i}$ and $\hat{\ell}_{i}$ are defined as in Lemma 6.3. Now we assume that the desired bound holds for $\ell_{1},\ldots,\ell_{i}$ and consider $\ell_{i+1}$. By the induction hypothesis, we have $\ell_{j}\geq d^{\varepsilon/15}$ for $j\leq i$. Furthermore, as we assume $d_{j}\geq\ell_{j}/8$ and $d$ is sufficiently large, we have $d_{j}\geq d^{\varepsilon/20}$. In particular, we may apply Lemmas 6.2 and 6.3. The following claim will assist with our proof:
@@
Note that as we assume $i<i^{\star}$, we have $\log\ell_{i}\geq\frac{\varepsilon}{15}\log d$ and $\log d_{i}\geq\frac{\varepsilon}{20}\log d$. Therefore, $\eta\gg\log^{-3}\ell_{i}$ and $\eta\gg\log^{-3}d_{i}$, and we can apply Lemmas 6.2 and 6.3.
@@
Acknowledgments
We are grateful to Luke Postle for bringing [27] to our attention. We also thank Eoin Hurley and Yuval Wigderson for helpful discussions.